A simple random sample of size n equals = 18 is drawn from a population that is normally distributed. The sample mean is found to be x overbar equals 54 and the sample standard deviation is found to be s equals = 19 Construct a 95​% confidence interval about the population mean. The 95​% confidence interval is ​( _____ ​, _____ ​). ​(Round to two decimal places as​ needed.)
Answer #1

561Utton Given thot 54 1 9 S 14 df=n-I = 18-1 957. Confidence level the t iS = 1- 95. 1 0.95 - o.05 o.025 2 2 tx/2 df to.o25,